# The Cayley transform on representations

### v1

@@ 4. The Cayley transform on a representation
Suppose $\mathbb{F}$ is either $\mathbb{R}$ or $\mathbb{C}$. Let $G$ be a Lie group and let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be a representation of $G$. We denote by $\mathfrak{g}$ the Lie algebra of $G$. Suppose $C:\mathfrak{gl}(\mathbb{V})\dasharrow\operatorname{GL}(\mathbb{V})$ is the classical Cayley transform defined by $C(u)=(1+u)/(1-u)$.
@@
We say that $\rho$ is a Cayley representation if there exist an open neighbourhood $V\subseteq\mathfrak{g}$ of 00 such that $C$ is defined on $d\rho(V)$ and $C(d\rho(V))\subseteq\rho(G)$. We also say that the Cayley representation is applicable to $\rho$.
Clearly, given a matrix Lie group $G\subseteq\operatorname{GL}_{n}(\mathbb{F})$, the natural inclusion $\iota:G\hookrightarrow\operatorname{GL}_{n}(\mathbb{F})$ is a Cayley representation if and only if $C$ is applicable to $G$ in the classical sense. The following basic properties are well-known for the classical Cayley transform.
@@
Let $G$ be a Lie group and let $\pi:G_{0}\to G$ be the universal cover of $G$. A representation $\rho:G\to\operatorname{GL}(\mathbb{V})$ is Cayley if and only if $\rho\circ\pi:G_{0}\to\operatorname{GL}(\mathbb{V})$ is Cayley. Moreover, if $G$ is real and $G^{\mathbb{C}}$ is its complexification, then $\rho$ is Cayley implies that $\operatorname{Ind}(\rho):G^{\mathbb{C}}\to\operatorname{GL}(\mathbb{V})$ is Cayley.
@@
The ’only if’ part follows immediately from the definition. We notice that $G_{0}$ and $G$ have the same Lie algebra $\mathfrak{g}$. If $\rho\circ\pi$ is Cayley, then we have $C(d(\rho\circ\pi)(V))\subseteq(\rho\circ\pi)(G_{0})$ from some open neighbourhood $V$ of $0\in\mathfrak{g}$. Since $\pi$ is a covering map, we have $d\pi=\operatorname{Id}_{\mathfrak{g}}$ and $\pi(G_{0})=G$. This implies $C(d\rho(V))\subseteq\rho(G)$. The ’moreover’ part follows immediately from the definition. ∎
@@
To characterize representations of $G$ to which the Cayley transform is applicable, we need the definition that follows.
@@
According to [13, Section 20.3], $\operatorname{Out}(\operatorname{Spin}_{8}(\mathbb{C}))$ acts on $\{\mathbb{C}^{n},S^{+},S^{-}\}$ as the permutation group $\mathfrak{S}_{3}$. Here $\mathbb{S}_{+},\mathbb{S}_{-}$ are the two spin representations. Since $\mathbb{C}^{n}$ is Cayley, Corollary 4.12 immediately implies that $\mathbb{S}_{+}$ and $\mathbb{S}_{-}$ are also Cayley representations.
@@ 5. The Cayley transform for semisimple Lie groups
Moreover, if $\mathfrak{g}$ is a non-compact real form, then $\mathfrak{g}$ admits a restricted root space decomposition [23, Proposition 6.40 and Equation (6.48b)]:
@@ 5.2. The Cayley transform on general representations
6. The Cayley transform for classical simple Lie groups
In this section, we consider the applicability of the Cayley transform to a representation $\rho:G\to\operatorname{GL}(\mathbb{V})$ of a classical semisimple Lie group $G$. We notice that the Lie algebra $\mathfrak{g}$ of $G$ is a simple Lie algebra, thus $d\rho:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V})$ is faithful unless $d\rho(\mathfrak{g})=\{0\}$. The latter case indicates that $\rho$ is the trivial representation. In the sequel, we suppose $\rho$ is non-trivial so that $d\rho$ is faithful.
@@ 6.1. Representations with Cayley configuration
We first classify Cayley representions of classical complex simple Lie groups. It is well-known [13, Theorem 21.11] that they are classified into four types: $A_{n},B_{n},C_{n}$ and $D_{n}$. Hence we may split our discussion correspondingly. For ease of reference, we record in Table 6.1 the information of the root system for each type. Here $\mathfrak{h}$ is the Cartan subalgebra of $\mathfrak{g}$, $\Lambda$ is the weight lattice, $\Phi^{+}$ is the set of positive roots, $\Delta^{+}$ is the set of simple roots, $C=\{\sum_{j=1}^{n}a_{j}L_{j}\}$ is the fundamental (closed) Weyl chamber, $\mathcal{W}$ is the Weyl group and $\delta\coloneqq\frac{1}{2}\sum_{j=1}^{n}L_{j}$. Moreover, in the last line of Table 6.1, we also list the standard representation $\mathbb{W}$ of each type together with its weights.
@@ 6.1.3. Type $C_{n}$ ($n\geq 3$)
Let $G$ be a complex simple Lie group of type $C_{n}$ ($n\geq 3$) and let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be a representation with Cayley configuration. Then we have $\mathbb{V}_{0}\neq 0$.
@@ 6.1.4. Type $D_{n}$ ($n\geq 4$)
Let $G$ be a complex simple Lie group of type $C_{n}$ ($n\geq 3$) and let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be a representation with Cayley configuration. Then we have $\mathbb{V}_{0}\neq 0$.
@@ 6.2. Classification
      $\mathfrak{g}\simeq\mathfrak{so}_{5}(\mathbb{C})$ and $\mathbb{V}$ is the spinor representation. Under the identification $\mathfrak{so}_{5}(\mathbb{C})\simeq\mathfrak{sp}_{4}(\mathbb{C})$, we have $\mathbb{V}\simeq\mathbb{C}^{4}$.
@@
Let $G$ be a compact real Lie group and let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be an irreducible representation of $G$. $\mathbb{F}=\mathbb{R}$. Suppose $G$ is simple. Let $\mathbb{V}=\oplus\mathbb{V}_{j}$ be the irreducible decomposition. The Cayley transform is applicable to $\rho$ if and only if for each $j$, $(\mathfrak{g},\mathbb{V}_{j})$ belongs to the following list
@@
    $\mathfrak{su}_{2}(\mathbb{C}),\mathfrak{su}_{2}(\mathbb{C})$. Under the identification $\mathfrak{su}_{2}(\mathbb{C})\simeq\mathfrak{so}_{3}(\mathbb{R})$, we have $\mathbb{V}\simeq\mathbb{C}^{3}$.
@@
    $(\mathfrak{so}_{5}(\mathbb{R}),\mathbb{S})$ where $\mathbb{S}$ is the spinor representation. Under the identification $\mathfrak{so}_{5}(\mathbb{R})\simeq\mathfrak{u}_{2}(\mathbb{H})$, we have $\mathbb{V}\simeq\mathbb{H}^{2}$. Here $\mathbb{H}$ denotes the quaternion algebra.

### v2

@@ 4. The Cayley transform on a representation
Suppose $\mathbb{F}$ is either $\mathbb{R}$ or $\mathbb{C}$. Let $G$ be a Lie group and let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be a representation of $G$. Here $\mathbb{V}$ is a finite-dimensional complex vector space. We denote by $\mathfrak{g}$ the Lie algebra of $G$. Assume $C:\mathfrak{gl}(\mathbb{V})\dasharrow\operatorname{GL}(\mathbb{V})$ is the classical Cayley transform defined by $C(u)=(1+u)/(1-u)$.
@@
We say that $\rho$ is a Cayley representation if there exist a neighbourhood $V_{1}\subseteq\mathfrak{g}$ of 00 and a neighbourhood $V_{2}\subseteq\mathrm{im}\rho$ [^1] The group $\mathrm{im}\rho$ is called the immersive Lie subgroup [47, Definition 4.2]. of the identity element such that $C$ is defined on $d\rho(V_{1})$ and $C(d\rho(V))\subseteq\rho(G)$. We also say that the Cayley transform is applicable to $\rho$.
Clearly, given a matrix Lie group $G\subseteq\operatorname{GL}_{n}(\mathbb{\mathbb{C}})$, the natural inclusion $\iota:G\hookrightarrow\operatorname{GL}_{n}(\mathbb{\mathbb{C}})$ is a Cayley representation if and only if $C$ is applicable to $G$ in the classical sense. The following basic properties are well-known for the classical Cayley transform.
@@
Let $G$ be a Lie group and let $\pi:G_{0}\to G$ be the universal cover of $G$. A representation $\rho:G\to\operatorname{GL}(\mathbb{V})$ is Cayley if and only if $\rho\circ\pi:G_{0}\to\operatorname{GL}(\mathbb{V})$ is Cayley.
@@
The ’only if’ part follows immediately from the definition. We notice that $G_{0}$ and $G$ have the same Lie algebra $\mathfrak{g}$. If $\rho\circ\pi$ is Cayley, then we have $C(d(\rho\circ\pi)(V))\subseteq(\rho\circ\pi)(G_{0})$ from some open neighbourhood $V$ of $0\in\mathfrak{g}$. Since $\pi$ is a covering map, we have $d\pi=\operatorname{Id}_{\mathfrak{g}}$ and $\pi(G_{0})=G$. This implies $C(d\rho(V))\subseteq\rho(G)$. ∎
We fix a norm on $\mathbb{V}$ and denote by $\lVert x\rVert$ the corresponding operator norm of $x\in\mathfrak{gl}(\mathbb{V})$.
@@
By the proof of Lemma 4.4, the bound $1/3$ in Lemma 4.4 is independent of the choice of the norm on $\mathbb{V}$. To characterize representations of $G$ to which the Cayley transform is applicable, we need the definition that follows.
@@
According to [15, Section 20.3], $\operatorname{Out}(\operatorname{Spin}_{8}(\mathbb{C}))$ acts on $\{\mathbb{C}^{n},S^{+},S^{-}\}$ as the permutation group $\mathfrak{S}_{3}$. Here $\mathbb{S}_{+},\mathbb{S}_{-}$ are the two spin representations. Since $\mathbb{C}^{n}$ is Cayley, Corollary 4.12 immediately implies that $\mathbb{S}_{+}$ and $\mathbb{S}_{-}$ are also Cayley representations.
@@ 5. The Cayley transform for semisimple Lie groups
Moreover, if $\mathfrak{g}$ is a non-compact real form, then $\mathfrak{g}$ admits a restricted root space decomposition [26, Proposition 6.40 and Equation (6.48b)]:
@@ 5.2. The Cayley transform on general representations
6. The Cayley transform for simple Lie groups
In this section, we consider the applicability of the Cayley transform to a representation $\rho:G\to\operatorname{GL}(\mathbb{V})$ of a classical simple Lie group $G$. In fact, one can prove that the Cayley transform is not applicable to exceptional simple Lie groups by similar calculations in Subsection 6.1. This is also a direct consequence of [31, Theorem 1.31].
We notice that the Lie algebra $\mathfrak{g}$ of $G$ is a simple Lie algebra, thus $d\rho:\mathfrak{g}\to\mathfrak{gl}(\mathbb{V})$ is faithful unless $d\rho(\mathfrak{g})=\{0\}$. The latter case indicates that $\rho$ is the trivial representation. In the sequel, we suppose $\rho$ is non-trivial so that $d\rho$ is faithful.
@@ 6.1. Representations with Cayley configuration
We first classify Cayley representions of classical complex simple Lie groups. It is well-known [15, Theorem 21.11] that they are classified into four types: $A_{n},B_{n},C_{n}$ and $D_{n}$. Hence we may split our discussion correspondingly. For ease of reference, we record in Table 6.1 the information of the root system for each type. Here $\mathfrak{h}$ is the Cartan subalgebra of $\mathfrak{g}$, $\Lambda$ is the weight lattice, $\Phi^{+}$ is the set of positive roots, $\Delta^{+}$ is the set of simple roots, $C=\{\sum_{j=1}^{n}a_{j}L_{j}\}$ is the fundamental (closed) Weyl chamber, $\mathcal{W}$ is the Weyl group and $\delta\coloneqq\frac{1}{2}\sum_{j=1}^{n}L_{j}$. Moreover, in the last line of Table 6.1, we also list the standard representation $\mathbb{W}$ of each type together with its weights.
@@ 6.1.3. Type $C_{n}$ ($n\geq 3$)
Let $G$ be a complex simple Lie group of type $C_{n}$ ($n\geq 3$) and let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be a representation with Cayley configuration. Then we have $\mathbb{V}_{0}=0$.
@@ 6.1.4. Type $D_{n}$ ($n\geq 4$)
Let $G$ be a complex simple Lie group of type $C_{n}$ ($n\geq 3$) and let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be a representation with Cayley configuration. Then we have $\mathbb{V}_{0}=0$.
@@ 6.2. Classification
      $\mathfrak{g}\simeq\mathbb{S})$ and $\mathbb{V}$ is the spin representation. Under the identification $\mathfrak{so}_{5}(\mathbb{C})\simeq\mathfrak{sp}_{4}(\mathbb{C})$, we have $\mathbb{V}\simeq\mathbb{C}^{4}$.
@@
Let $G$ be a compact real simple Lie group and let $\rho:G\to\operatorname{GL}(\mathbb{V})$ be an irreducible representation of $G$. The Cayley transform is applicable to $\rho$ if and only if $(\mathfrak{g},\mathbb{V})$ belongs to the following list
@@
    $(\mathfrak{su}_{2}(\mathbb{C}),\mathfrak{sl}_{2}(\mathbb{C}))$, where $\mathfrak{sl}_{2}(\mathbb{C})$ is the complexification of the adjoint representation of $\mathfrak{su}_{2}(\mathbb{C})$. Under the identification $\mathfrak{su}_{2}(\mathbb{C})\simeq\mathfrak{so}_{3}(\mathbb{R})$, we have $\mathbb{V}\simeq\mathbb{C}^{3}$.
@@
    $(\mathfrak{so}_{5}(\mathbb{R}),\mathbb{S})$. $\mathbb{S}$ is the spinor representation. Under the identification $\mathfrak{so}_{5}(\mathbb{R})\simeq\mathfrak{u}_{2}(\mathbb{H})$, we have $\mathbb{V}\simeq\mathbb{H}^{2}$. Here $\mathbb{H}$ denotes the quaternion algebra.